# Cut-based Conflict Analysis in Mixed Integer Programming

### v1

@@ 1 Introduction
Although the use of conflict analysis in so-called conflict-driven clause learning (CDCL) SAT solvers has been hugely successful, one drawback is that from a mathematical point of view the resolution proof system on which it is based is quite weak, and is known to require proofs of exponential length even for simple combinatorial principles (Haken 1985, Urquhart 1987). The requirement to encode the input in conjunctive normal form (CNF) as a collection of disjunctive clauses incurs a further loss in expressive power. It has therefore been studied how to lift SAT-based conflict-driven methods to richer input formats such as (linear) pseudo-Boolean (PB) constraints, which translated from SAT to MIP is just another name for 001111 integer linear programs with integer coefficients. Crucially, here it turns out that the two equivalent ways of describing SAT conflict analysis discussed above generalize in different directions.
Many pseudo-Boolean solvers work by translating the input to CNF, possibly by introducing auxiliary variables, and then perform search and conflict analysis on this representation (Eén and Sörensson 2006, Martins et al. 2014, Joshi et al. 2015, Sakai and Nabeshima 2015). Except for any auxiliary variables, this works the same as the graph-based conflict analysis hitherto used in MIP solvers Achterberg (2007a). Another approach, however, is to adopt the derivation-based view, but to perform the conflict analysis derivation in a proof system adapted to 001111 linear inequalities. The natural candidate for such a proof system is cutting planes (Cook et al. 1987), which has been extensively studied in the area of computational complexity theory. It is a priori not obvious what it would mean to perform conflict analysis in the cutting planes proof system, but methods to do so have been designed in (Dixon and Ginsberg 2002, Chai and Kuehlmann 2005a, Sheini and Sakallah 2006) and are currently used in the pseudo-Boolean solvers Sat4j (Le Berre and Parrain 2010) and RoundingSat (Elffers and Nordström 2018). As a theoretical method of reasoning, such cut-based conflict analysis turns out to be exponentially more powerful than the graph-based conflict analysis yielding resolution proofs.
@@
We remark for completeness that there is a third way of learning from local infeasibilities in MIP called dual-proof analysis (Witzig et al. 2019). It is conceptually different from both graph- and cut-based conflict analysis, in that (i) it can only be applied when the infeasibility has been detected via an infeasible LP relaxation, not by propagation, and (ii) the learned dual-proof constraint does not depend on the history of bound changes but is constructed in a single step. A notable commonality with cut-based conflict analysis is that dual-proof constraints and cut-based learned constraints are derived by aggregating a subset of the original input constraints, and so both methods operate syntactically on linear constraints. One important difference is that in cut-based conflict analysis, the set of constraints to aggregate and the multipliers to use are computed based on the propagations that have been implied, and also that cuts can be applied to the individual constraints before aggregation. In dual-proof analysis, no cuts are needed since the LP relaxation is already infeasible, and the multipliers are obtained from the dual solution for the LP relaxation. Pseudo-Boolean solving has also been combined with LP solving of relaxed versions of the input problem, and there dual-proof analysis has been combined with pseudo-Boolean conflict analysis (Devriendt et al. 2021).
@@ 1.1 Questions Studied in This Work and Our Contributions
Our focus in this work is on how cut-based conflict analysis as described above can be integrated in MIP solvers. A first step was taken in (Mexi et al. 2023) by showing how to apply this method for pure binary programs. In this paper, we go much further by applying cut-based conflict analysis in general MIP solving. We design and implement a new conflict analysis method that is guaranteed to work not only for 001111 integer linear programs but also in the presence of continuous variables. We also extend this approach to general integer variables, but here the method is not always guaranteed to work when there are integer variable bound changes in the proper interior of the variable’s domain.
@@
We have implemented all of these reduction and conflict analysis methods in the open-source MIP solver SCIP, and present the results of a thorough computational study on MIP instances from MIPLIB 2017 with quantitative data for the performance of different conflict analysis algorithms and also qualitative data about properties of the learned constraints. Our experiments indicate that our proposed conflict analysis algorithm generates useful conflict constraints and improves the default performance of the MIP solver SCIP not only in terms of running time and number of nodes in the search tree, but also in terms of the number of instances solved.
@@ 1.2 Organization of This Paper
After a review of preliminaries in Section 2, we formalize our MIP interpretation of pseudo-Boolean cut-based conflict analysis in Section 3. Section 4 presents a collection of different reduction algorithms, and Section 5 studies dominance relationships between them. We extend cut-based conflict analysis to mixed binary programs in Section 6, and discuss general MIP problems with integer variables in Section 7. The results from our computational evaluation are presented in Section 8, after which we provide some concluding remarks and discuss future research directions in Section 9.
@@ 2 Preliminaries and Notation
Depending on the set $\mathcal{I}$ and the bound vectors $\ell$ and $u$, we distinguish the following special cases of MIPs. We call (1) a mixed binary program (MBP) if $\ell_{j}=0$ and $u_{j}=1$ for all $j\in\mathcal{I}$. If additionally =$\mathcal{I}=\{1,\dots,n\}$, then we call (1) a pure binary program (BP).
@@ 2.1 States and Activities
For a generic constraint $C:\sum_{j}a_{j}x_{j}\geq b$, we denote the maximal activity of $C$ under state $\rho$ as
$\textnormal{act}^{\text{max}}({C,\rho}):=\sum_{j}\max\{a_{j}\cdot{\ell_{j}^{\rho}},\,a_{j}\cdot u_{j}^{\rho}\}.$
Similarly, we define the minimal activity of $C$ under $\rho$ as
$\textnormal{act}^{\text{min}}({C,\rho}):=\sum_{j}\min\{a_{j}\cdot{\ell_{j}^{\rho}},\,a_{j}\cdot u_{j}^{\rho}\}.$
We call the constraint $C$ infeasible under some state if its maximal activity is less than the right-hand side, i.e., if the respective inequality cannot hold for any solution within the local bounds. Otherwise, we call $C$ feasible. We are particularly interested in the fact whether the bound changes on a variable affect the current maximal activity of a constraint. We call a variable $x_{j}$ relaxable for constraint $C$ under state $\rho$ if $\textnormal{act}^{\text{max}}({C,\rho})$ remains unchanged when we replace its local bounds ${\ell_{j}^{\rho}}$ and $u_{j}^{\rho}$ by its global bounds $\ell_{j}$ and $u_{j}$, respectively. By definition, this is the case if and only if $a_{j}\not=0$ or $a_{j}>0\land u_{j}^{\rho}=u_{j}$ or $a_{j}<0\land{\ell_{j}^{\rho}}=\ell_{j}$. Otherwise, the variable is called non-relaxable. Note that these terms provide a generalization from the PB to the MIP setting. In the literature on pseudo-Boolean optimization, where there are only binary variables, relaxable and non-relaxable variables correspond to non-falsified and falsified literals, respectively.
@@ 2.2 Propagation of Linear Constraints
Beyond identifying trivial infeasibility, the maximal activity of a constraint under some state $\rho$ can also be used to identify whether variable bounds can be tightened. A commonly used propagator for linear constraints is the bound strengthening technique going back to Brearley et al. (1975). If $\textnormal{act}^{\text{max}}({C,\rho})<\infty$, then we can rewrite the constraint $C$ as constraint $C$ as
@@
which can be used to tighten the lower bound of $x_{r}$ if ~$\tilde{\ell}_{r}>{\ell_{r}^{\rho}}$. In a similar fashion, if $a_{r}<0$, we can deduce the upper bound $x_{r}\leq\tilde{u}_{r}:={\ell_{r}^{\rho}}+\frac{b-\textnormal{act}^{\text{max}}({C,\rho})}{a_{r}}$ if $\tilde{u}_{r}<u_{r}^{\rho}$. For completeness, note that if $\textnormal{act}^{\text{max}}({C,\rho})=\infty$ because $a_{r}>0\land u_{r}^{\rho}=\infty$ or $a_{r}<0\land{\ell_{r}^{\rho}}=-\infty$, but $\max_{x\in[{\ell^{\rho}},u^{\rho}]}\sum_{j\neq r}a_{j}x_{j}<\infty$, then (2) can be used directly to derive valid bounds.
For integer variables, i.e., for $r\in\mathcal{I}$, the derived bounds can be rounded to ~$\lceil\tilde{\ell}_{r}\rceil$ and $\lfloor\tilde{u}_{r}\rfloor$, respectively. As will become clear in Section 3, we are particularly interested in propagations where this second step is redundant because either the variable $x_{r}$ is continuous or the propagated bounds are already integral, i.e., ~$\tilde{\ell}_{r}\in\mathbb{Z}$ and $\tilde{u}_{r}\in\mathbb{Z}$, respectively. We refer to such propagations as tight.
@@ 2.3 Coefficient Tightening and Cutting Planes
Note that the above procedure can be applied only for non-redundant constraints, i.e., for constraints where $\textnormal{act}^{\text{min}}({C})<b$. If all variables in $C$ are binary with non-negative coefficients then coefficient tightening is known as saturation (Chai and Kuehlmann 2005a).
Definition 2 (Saturation Cut).
Let $C:\sum_{j\in J}a_{j}x_{j}\geq b$ with $x\in\{0,1\}^{J}$, $J\subseteq\mathcal{I}$, and $a_{j}\geq~{}0,j\in~{}J$. The saturation cut of $C$ is given by the constraint
$\sum_{j\in J}\min\{a_{j},b\}x_{j}\geq b.$
@@
Finally, we recall the Mixed Integer Rounding (MIR) cut Marchand and Wolsey (2001), which is a generalization of Gomory’s mixed integer cuts Gomory (1960).
@@
$\sum_{\begin{subarray}{c}j\in K:\\ a_{j}>0\end{subarray}}\frac{a_{j}}{f(b)}x_{j}+\sum_{j\in J}\left(\lfloor a_{j}\rfloor+\min\left\{1,\frac{f(a_{j})}{f(b)}\right\}\right)x_{j}\geq\left\lceil b\right\rceil.$ (6)
@@ 2.4 Weakening and Complementation
Two important operations on constraints that are used in later sections are weakening and complementation of variables. Weakening a variable $x_{s}$ in a constraint $C:\,\sum_{j}a_{j}x_{j}\geq b$ is defined as
$\text{weaken}(C,\,x_{s}):=\sum_{j\neq s}a_{j}x_{j}\geq b-\text{max}\{a_{s}u_{s},a_{s}\ell_{s}\}.$
Weakening is a valid operation since it simply adds a multiple of the globally valid bound constraints $x_{j}\leq u_{j}\Leftrightarrow-x_{j}\geq-u_{j}$, or $x_{j}\geq\ell_{j}$ to $C$. Weakening entails a loss of information. However, as we will see, it is a necessary operation in some reduction algorithms. Note that whenever weakening is applied on relaxable variables at the current state, it does not change the bound propagations of the remaining variables in the constraint.
The second operation that we use in the reduction is complementation of variables. Complementation is a valid operation that does not entail a loss of information since it simply replaces a variable $x_{s}$ by its complement $\overline{x}_{s}=u_{s}-x_{s}$. Complementing a variable $x_{s}$ in the general constraint $C$ from above is defined as
$\text{complement}(C,\,x_{s}):=\sum_{j\neq s}a_{j}x_{j}-a_{s}\overline{x}_{s}\geq b-a_{s}u_{s}.$
@@ 3 The Simple Case: Conflict Analysis under Tight Propagations
Conflict analysis is applicable whenever a locally infeasible constraint is detected at a node of the branch-and-bound search. In MIP solvers, this constraint may be detected directly during propagation or as trivially infeasible aggregation of constraints resulting from an infeasible LP relaxation (Achterberg 2007a). This can occur after many propagations on the current node, and the main goal of conflict analysis is to obtain a constraint that would have identified the infeasibility earlier. Such a constraint would have propagated in an earlier node, preventing us from visiting this infeasible node in the first place. [^1] This is also called non-chronological backtracking in the SAT literature Marques-Silva and Sakallah (1996). The technique for constructing such a constraint is as follows.
Given the constraint that found the infeasibility, called conflict constraint, and the state history, we can find the last constraint that propagated a variable at a state $\rho$ which contributed to the infeasibility of the conflict constraint, called reason constraint. The conflict and reason constraints form a system of two constraints that cannot be satisfied simultaneously at a state $\tilde{\rho}$ before $\rho$. Now, the goal is to obtain a single globally valid constraint that is infeasible under $\tilde{\rho}$. If we have a general recipe to achieve this, this process can be repeated, using the learned constraint as the new conflict constraint. A common criterion is to iterate this step until a so-called first unique implication point (FUIP) is reached, i.e., until the learned constraint would propagate some variable at a state in a previous decision level. Note that the learned constraint may also be infeasible under the global bounds, in which case we have proven that the problem is globally infeasible. To summarize, the fundamental problem of conflict analysis is: Given two rows $C_{\textnormal{confl}}$ and $C_{\textnormal{reason}}$, a local domain $L$, and a global domain $G$, such that the system $\{C_{\textnormal{confl}},C_{\textnormal{reason}},L\}$ is infeasible, find a constraint $C_{\textnormal{learn}}$ that is valid for $\{C_{\textnormal{confl}},C_{\textnormal{reason}},G\}$ such that $\{C_{\textnormal{learn}},L\}$ is infeasible.
@@
We can interpret the above technique as trying to eliminate $x_{r}$ from the system $\{C_{\textnormal{confl}},C_{\textnormal{reason}},L\}$ by Fourier-Motzkin elimination. Indeed, the system is still infeasible even when we relax the domain of $x_{r}$ to (,)($\infty$,$\infty$)( - ∞ , ∞ ). Then, by Fourier-Motzkin elimination Williams (1976), projecting out $x_{r}$ yields exactly (7) Note that this immediately proves the infeasibility of (7) in $L$, since the projection of the empty set is the empty set.
@@
$-\beta+\max_{x\in[{\ell^{\rho}},u^{\rho}]}\sum_{j\neq r}a^{\prime}_{j}x_{j}<b^{\prime}$
@@
$=b-\beta+\max_{x\in[{\ell^{\rho}},u^{\rho}]}\sum_{j\neq r}a^{\prime}_{j}x_{j}$
$<b+b^{\prime},$
which shows that $C_{\text{res}}$ is infeasible in the state $\rho$.
@@
Non-tight propagation can occur if and only if the optimum of (9) over the current state is attained at a non-integer vertex, as illustrated in Figure 1 for the reason constraint of Example 1. By contrast, for SAT constraints the feasible region defined by the reason constraint and the global domain does not contain non-integer vertices.
@@
Figure 1: Inequality $x_{1}+x_{2}+2x_{3}\geq 2$ on the face of the polytope with $x_{1}=0$. Here the fractional vertex is $(0,1,0.5)$.
This geometric perspective also shows a clear path forward to make conflict analysis work: by strengthening the propagating reason constraint $C_{\textnormal{reason}}$ in order to cut off the non-integer vertex from the feasible region $C_{\textnormal{reason}}\cap G$. Note that for pure binary programs the reason is a knapsack constraint, and hence it is sufficient to study cuts for the knapsack polytope Hojny et al. (2020). In the PB literature, the techniques to achieve such cuts are called reduction techniques. In the next section, we first present the cut-based conflict analysis algorithm and then proceed to discuss various reduction techniques.

### v2

@@ 1 Introduction
Although the use of conflict analysis in so-called conflict-driven clause learning (CDCL) SAT solvers has been hugely successful, one drawback is that from a mathematical point of view the resolution proof system on which it is based is quite weak, and is known to require proofs of exponential length even for simple combinatorial principles (Haken 1985, Urquhart 1987). The requirement to encode the input in conjunctive normal form (CNF) as a collection of disjunctive clauses incurs a further loss in expressive power. It has therefore been studied how to lift SAT-based conflict-driven methods to richer input formats such as (linear) pseudo-Boolean (PB) constraints, which translated from SAT to MIP is just another name for 011 integer linear programs with integer coefficients. Crucially, here it turns out that the two equivalent ways of describing SAT conflict analysis discussed above generalize in different directions.
Many pseudo-Boolean solvers work by translating the input to CNF, possibly by introducing auxiliary variables, and then perform search and conflict analysis on this representation (Eén and Sörensson 2006, Martins et al. 2014, Joshi et al. 2015, Sakai and Nabeshima 2015). Except for any auxiliary variables, this works the same as the graph-based conflict analysis hitherto used in MIP solvers Achterberg (2007a). Another approach, however, is to adopt the derivation-based view, but to perform the conflict analysis derivation in a proof system adapted to 011 linear inequalities. The natural candidate for such a proof system is cutting planes (Cook et al. 1987), which has been extensively studied in the area of computational complexity theory. It is a priori not obvious what it would mean to perform conflict analysis in the cutting planes proof system, but methods to do so have been designed in (Dixon and Ginsberg 2002, Chai and Kuehlmann 2005, Sheini and Sakallah 2006) and are currently used in the pseudo-Boolean solvers Sat4j (Le Berre and Parrain 2010) and RoundingSat (Elffers and Nordström 2018). As a theoretical method of reasoning, such cut-based conflict analysis turns out to be exponentially more powerful than the graph-based conflict analysis yielding resolution proofs.
@@
We remark for completeness that there is a third way of learning from local infeasibilities in MIP called dual-proof analysis (Witzig et al. 2021). It is conceptually different from both graph- and cut-based conflict analysis, in that (i) it can only be applied when the infeasibility has been detected via an infeasible LP relaxation, not by propagation, and (ii) the learned dual-proof constraint does not depend on the history of bound changes but is constructed in a single step. A notable commonality with cut-based conflict analysis is that dual-proof constraints and cut-based learned constraints are derived by aggregating a subset of the original input constraints, and so both methods operate syntactically on linear constraints. One important difference is that in cut-based conflict analysis, the set of constraints to aggregate and the multipliers to use are computed based on the propagations that have been implied, and also that cuts can be applied to the individual constraints before aggregation. In dual-proof analysis, no cuts are needed since the LP relaxation is already infeasible, and the multipliers are obtained from the dual solution for the LP relaxation. Pseudo-Boolean solving has also been combined with LP solving of relaxed versions of the input problem, and there dual-proof analysis has been combined with pseudo-Boolean conflict analysis (Devriendt et al. 2021).
@@ 1.1 Questions Studied in This Work and Our Contributions
Our focus in this work is on how cut-based conflict analysis as described above can be integrated in MIP solvers. A first step was taken in (Mexi et al. 2023) by showing how to apply this method for pure binary programs. In this paper, we go much further by applying cut-based conflict analysis in general MIP solving. We design and implement a new conflict analysis method that is guaranteed to work not only for 011 integer linear programs but also in the presence of continuous variables. We also extend this approach to general integer variables, but here the method is not always guaranteed to work when there are integer variable bound changes in the proper interior of the variable’s domain.
@@
We have implemented all of these reduction and conflict analysis methods in the MIP solver SCIP, and present the results of our computational study on MIP instances from MIPLIB 2017 with quantitative data for the performance of different conflict analysis algorithms and also qualitative data about properties of the learned constraints. Our experiments indicate that our proposed conflict analysis algorithm generates useful conflict constraints and improves the default performance of the MIP solver SCIP not only in terms of running time and number of nodes in the search tree, but also in terms of the number of instances solved.
@@ 1.2 Organization of This Paper
After a review of preliminaries in Section˜2, we formalize our MIP interpretation of pseudo-Boolean cut-based conflict analysis in Section˜3. Section˜4 presents a collection of different reduction algorithms, and Section˜5 studies dominance relationships between them. We extend cut-based conflict analysis to mixed binary programs in Section˜6, and discuss general MIP problems with integer variables in Section˜7. The results from our computational evaluation are presented in Section˜8, after which we provide some concluding remarks and discuss future research directions in Section˜9.
@@ 2 Preliminaries and Notation
Depending on the set $\mathcal{I}$ and the bound vectors $\ell$ and $u$, we distinguish the following special cases of MIPs. We call (1) a mixed binary program (MBP) if $\ell_{j}=0$ and $u_{j}=1$ for all $j\in\mathcal{I}$. If additionally $\mathcal{I}=\{1,\dots,n\}$, then we call (1) a pure binary program (BP).
@@ 2.1 States and Activities
For a generic constraint $C:\sum_{j}a_{j}x_{j}\geq b$, we denote the maximal activity of $C$ under state $\rho$ as $\textnormal{act}^{\text{max}}({C,\rho}):=\sum_{j}\max\{a_{j}\cdot{\ell_{j}^{\rho}},\,a_{j}\cdot u_{j}^{\rho}\}.$ Similarly, we define the minimal activity of $C$ under $\rho$ as $\textnormal{act}^{\text{min}}({C,\rho}):=\sum_{j}\min\{a_{j}\cdot{\ell_{j}^{\rho}},\,a_{j}\cdot u_{j}^{\rho}\}.$ We call the constraint $C$ infeasible under some state if its maximal activity is less than the right-hand side, i.e., if the respective inequality cannot hold for any solution within the local bounds. Otherwise, we call $C$ feasible. We are particularly interested in the fact whether the bound changes on a variable affect the current maximal activity of a constraint. We call a variable $x_{j}$ relaxable for constraint $C$ under state $\rho$ if $\textnormal{act}^{\text{max}}({C,\rho})$ remains unchanged when we replace its local bounds $\ell_{j}^{\rho}$} and $u_{j}^{\rho}$ by its global bounds $\ell_{j}$ and $u_{j}$, respectively. By definition, this is the case if and only if $a_{j}=0$ or $a_{j}>0\land u_{j}^{\rho}=u_{j}$ or $a_{j}<0\land{\ell_{j}^{\rho}}=\ell_{j}$. Otherwise, the variable is called non-relaxable. Note that these terms provide a generalization from the PB to the MIP setting. In the literature on pseudo-Boolean optimization, where there are only binary variables, relaxable and non-relaxable variables correspond to non-falsified and falsified literals, respectively.
@@ 2.2 Propagation of Linear Constraints
Beyond identifying trivial infeasibility, the maximal activity of a constraint under some state $\rho$ can also be used to identify whether variable bounds can be tightened. A commonly used propagator for linear constraints is the bound strengthening technique going back to Brearley et al. (1975). If $\textnormal{act}^{\text{max}}({C,\rho})<\infty$, then we can rewrite constraint $C$ as
@@
which can be used to tighten the lower bound of $x_{r}$ if $\tilde{\ell}_{r}>{\ell_{r}^{\rho}}$. In a similar fashion, if $a_{r}<0$, we can deduce the upper bound $x_{r}\leq\tilde{u}_{r}:={\ell_{r}^{\rho}}+\frac{b-\textnormal{act}^{\text{max}}({C,\rho})}{a_{r}}$ if $\tilde{u}_{r}<u_{r}^{\rho}$. For completeness, note that if $\textnormal{act}^{\text{max}}({C,\rho})=\infty$ because $a_{r}>0\land u_{r}^{\rho}=\infty$ or $a_{r}<0\land{\ell_{r}^{\rho}}=$-$\infty$, but $\max_{x\in[{\ell^{\rho}},u^{\rho}]}\sum_{j\neq r}a_{j}x_{j}<\infty$, then (2) can be used directly to derive valid bounds.
For integer variables, i.e., for $r\in\mathcal{I}$, the derived bounds can be rounded to $\lceil\tilde{\ell}_{r}\rceil$ and $\lfloor\tilde{u}_{r}\rfloor$, respectively. As will become clear in Section˜3, we are particularly interested in propagations where this second step is redundant because either the variable $x_{r}$ is continuous or the propagated bounds are already integral, i.e., $\tilde{\ell}_{r}\in\mathbb{Z}$ and $\tilde{u}_{r}\in\mathbb{Z}$, respectively. We refer to such propagations as tight.
@@ 2.3 Coefficient Tightening and Cutting Planes
Next, we recall the Mixed Integer Rounding (MIR) cut Marchand and Wolsey (2001).
@@
$\sum_{\begin{subarray}{c}j\in K:\\ a_{j}>0\end{subarray}}\frac{a_{j}}{f(b)}x_{j}+\sum_{j\in J}\left(\lfloor a_{j}\rfloor+\min\left\{1,\frac{f(a_{j})}{f(b)}\right\}\right)x_{j}\geq\left\lceil b\right\rceil,$ (6)
where $f(r)=r-\lfloor r\rfloor$ is the fractional part of $r$.
@@ 2.4 Weakening and Complementation
Two operations on constraints that are used in later sections are weakening and complementation of variables. Weakening is a valid operation since it simply adds a multiple of the globally valid bound constraints $x_{j}\leq u_{j}\Leftrightarrow-x_{j}\geq-u_{j}$, or $x_{j}\geq\ell_{j}$ to $C$. Weakening a variable $x_{s}$ in a constraint $C:\,\sum_{j}a_{j}x_{j}\geq b$ is defined as: $\text{weaken}(C,\,x_{s}):=\sum_{j\neq s}a_{j}x_{j}\geq b-\text{max}\{a_{s}u_{s},a_{s}\ell_{s}\}.$ Weakening entails a loss of information. However, as we will see, it is a necessary operation in some reduction algorithms. Note that whenever weakening is applied on relaxable variables at the current state, it does not change the bound propagations of the remaining variables in the constraint.
The second operation that we use in the reduction is complementation of variables. Complementation is a valid operation that does not entail a loss of information since it simply replaces a variable $x_{s}$ by its complement $\overline{x}_{s}=u_{s}-x_{s}$. Complementing a variable $x_{s}$ in the general constraint $C$ from above is defined as: $\text{complement}(C,\,x_{s}):=\sum_{j\neq s}a_{j}x_{j}-a_{s}\overline{x}_{s}\geq b-a_{s}u_{s}.$
@@ 3 The Simple Case: Conflict Analysis under Tight Propagations
Conflict analysis is applicable whenever a locally infeasible constraint is detected at a node of the branch-and-bound search. In MIP solvers, this constraint may be detected directly during propagation or as trivially infeasible aggregation of constraints resulting from an infeasible LP relaxation (Achterberg 2007a). This can occur after many propagations on the current node, and the main goal of conflict analysis is to obtain a constraint that would have identified the infeasibility earlier. Such a constraint would have propagated in an earlier node, preventing us from visiting this infeasible node in the first place.The technique for constructing such a constraint is as follows.
Given the constraint that found the infeasibility, called conflict constraint, and the state history, we can find the last constraint that propagated a variable at a state $\rho$ which contributed to the infeasibility of the conflict constraint, called reason constraint. The conflict and reason constraints form a system of two constraints that cannot be satisfied simultaneously at a state $\tilde{\rho}$ before $\rho$. Now, the goal is to obtain a single globally valid constraint that is infeasible under $\tilde{\rho}$. If we have a general recipe to achieve this, this process can be repeated, using the learned constraint as the new conflict constraint. A common criterion is to iterate this step until a so-called first unique implication point (FUIP) is reached, i.e., until the learned constraint would propagate some variable at a state in the previous decision level. Note that the learned constraint may also be infeasible under the global bounds, in which case we have proven that the problem is globally infeasible. To summarize, the fundamental problem of conflict analysis is: Given two constraints $C_{\textnormal{confl}}$ and $C_{\textnormal{reason}}$, a local domain $L$, and a global domain $G$, such that the system $\{C_{\textnormal{confl}},C_{\textnormal{reason}},L\}$ is infeasible, find a constraint $C_{\textnormal{learn}}$ that is valid for $\{C_{\textnormal{confl}},C_{\textnormal{reason}},G\}$ such that $\{C_{\textnormal{learn}},L\}$ is infeasible.
@@
We can interpret the above technique as trying to eliminate $x_{r}$ from the system $\{C_{\textnormal{confl}},C_{\textnormal{reason}},L\}$ by Fourier-Motzkin elimination. Indeed, the system is still infeasible even when we relax the domain of $x_{r}$ to $(-\infty,\infty)$. Then, by Fourier-Motzkin elimination Williams (1976), projecting out $x_{r}$ yields exactly (7) Note that this immediately proves the infeasibility of (7) in $L$, since the projection of the empty set is the empty set.
@@
$\beta+\max_{x\in[{\ell^{\rho}},u^{\rho}]}\sum_{j\neq r}a^{\prime}_{j}x_{j}<b^{\prime}$
@@
$=b-\beta+\max_{x\in[{\ell^{\rho}},u^{\rho}]}\sum_{j\neq r}a^{\prime}_{j}x_{j}<b+b^{\prime},$
which shows that $C_{\text{res}}$ is infeasible in the state $\rho$. ∎
@@
Non-tight propagation can occur if and only if the optimum of (9) over the current state is attained at a non-integer vertex, as illustrated in Figure˜1 for the reason constraint of Example˜1. By contrast, for SAT constraints the feasible region defined by the reason constraint and the global domain does not contain non-integer vertices. This geometric perspective also shows a clear path forward to make conflict analysis work: by strengthening the propagating reason constraint $C_{\textnormal{reason}}$ in order to cut off the non-integer vertex from the feasible region $C_{\textnormal{reason}}\cap G$. Note that for pure binary programs the reason is a knapsack constraint, and hence it is sufficient to study cuts for the knapsack polytope Hojny et al. (2020). In the PB literature, the techniques to achieve such cuts are called reduction techniques. In the next section, we first present the cut-based conflict analysis algorithm and then proceed to discuss various reduction techniques.
@@
Figure 1: Inequality $x_{1}+x_{2}+2x_{3}\geq 2$ on the face of the polytope with $x_{1}=0$. The fractional vertex is $(0,1,0.5)$.